OpenStudy (anonymous):

The formula F = ma shows the relationship between force, mass, and acceleration. Solve this formula for a. a = F + m a = F – m a = Fm a = F divided by m

OpenStudy (anonymous):

@justonelastgoojudbye

OpenStudy (kirstin012596):

a=f/m

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

one more Bonnie has a container in the shape of a rectangular pyramid. The formula for the surface area of the enclosed space is S = lw + 0.5Ph. Solve for P. P = S – lw – 0.5h P = S + lw + 0.5h P = The quantity S minus l times w all divided by 0.5 times h P = S divided by the quantity l times w plus 0.5 times h

OpenStudy (kirstin012596):

yes, its just like a teenage problem. F=ma so treat a like x in a teenage problem, try to get it alone. you must divide f by m in order to move it to the other side. you pretty much have to do the opposite of what m is doing to a. m is being multiplied by a so you have to divide by m. if it was f=m+a you would do f-m=a. does that make sense?

OpenStudy (anonymous):

so is it a,b,c,or d

OpenStudy (kirstin012596):

the first one is D, and the second question I believe is C.

OpenStudy (anonymous):

ok thank you

OpenStudy (kirstin012596):

no problem